The following information pertains to Parsons Co.:

Preferred stock, cumulative:
    Par value per share $100
    Dividend rate 8%
    Shares outstanding 11,000
    Dividends in arrears none
Common stock:
    Par value per share $10
    Shares issued 125,000
    Dividends paid per share $2.10
    Market price per share $47.00
Additional paid-in capital $490,000
Unappropriated retained earnings (after closing) $250,000
Retained earnings appropriated for contingencies $300,000
Common treasury stock:
    Number of shares 11,000
    Total cost $250,000
Net income $633,000


Compute (assume no changes in balances during the past year): (Round per share and ratios to 2 decimal places, e.g. $15.25 or 15.25%.)

(a) Total amount of stockholders' equity in the balance sheet $
(b) Earnings per share of common stock $ per share
(c) Book value per share of common stock $ per share
(d) Payout ratio of common stock %
(e) Return on common stock equity %
Answer #1

Ans:

a) Total amount of stockholder's equity in the balance sheet

Preferred stock (11000*100)

1100,000

common stock (125000*10)

1250,000

Additional paid in capital

490,000

inappropriate Retained Earnings

250,000

Retained Earnings

300,000

Treasury stock

-250000

Total amount of stockholder's equity

3140,000

b) Earnings per share of common stock

Net income -dividend on preferred stock/common stock-common treasury stock

preferred stock dividend = (11000*100)*8% = 88000

Earning per share = (633000-88000)/(125000-11000)

earning per share = 545000/114000 = 4.78 per share

c) Book value per share of common stock

stockholder's equity-value of preferred stock/common stock-common treasury stock

Book value per share = (3140,000-1100000)/(125000-11000)

Book value per share = 17.8947=17.90

d) Payout ratio of common stock = Dividend per share/Earning per share

Payout ratio of common stock = 2.10/4.78

Payout ratio of common stock = 43.93%

e) Return on common stock equity

(633,000-88000)/( (3140,000-1100000)= 545000/2040,000=26.72%